# An Empirical Analysis on the Use and Reporting of National Security Letters

### 07 Mar 2024 v2

@@ Abstract
In this paper, we attempt to improve this state of affairs by consolidating the available data around NSLs from various sources and analyzing it. Our data collection allows us to answer two questions: (1) what can the public effectively learn from the reported data and does this information suffice to assess the NSL usage? and (2) how accessible is this data collection? On the one hand, we show that after consolidating and processing data, longitudinal trends in the usage of NSLs can be observed. For instance, we find a significant increase in NSL requests for non-US persons and that the policy reforms to decrease the mandated nondisclosure period appear to be effective. The observed trends suggest that the current transparency mechanisms are viable safeguards against the excessive use of NSLs. On the other hand, we find that aggregating and normalizing the data is a challenging task, and one of our main contributions is producing a normalized and accessible data set and making it available for future research. The various data sources pose major challenges as they all lack a standardized and machine-readable format. Releasing transparency reports and NSLs, after their gag order was lifted, is optional for companies and only done consistently by a few. While the government data is comparatively straightforward to find, as the publishing entities are known, processing them is challenging. Indeed, the format of the reported data changed repeatedly and the publication as scanned PDFs with continuous text requires manual reviewing, parsing, and validating of the data. We even find inconsistencies within and across data sources. Overall, the laborious data collection process hinders external and internal auditing efforts and demonstrates the need for a unified and more usable dataset for NSLs. We advocate for publishing NSL statistics in a standardized and easily processable format to facilitate third-party data analysis that ensures that Congress and the public can satisfy their role of overseeing the use of NSLs by government agencies.
@@ 3.3 Company NSLs
Figure 4: This graph shows how many years it took to publish NSLs after they were issued. Most letters were only published after the USA FREEDOM Act added procedures to challenge gag orders. The gag order time appears to decrease. However, the distribution of 99.7% of the NSLs is unknown. Moreover, the gag order time is biased towards shorter times, as letters from recent years with longer times have not been published yet.
However, two important caveats need to be mentioned in connection with Figure 4. First, we only see 0.3% of all NSLs, and it is unknown whether they are a representative sample for the distribution of the remaining 99.7% of NSLs. Second, data since 2013 is increasingly skewed towards shorter gag order times, as any letter that is in the “future dates” area will only be published in the future. For instance, letters issued in 2022 with a gag order time of four years will only be published in 2026.

### 11 Apr 2024 v3

@@ Abstract
In this paper, we attempt to improve this state of affairs by consolidating the available data around NSLs from various sources and analyzing it. Our data collection allows us to answer two questions: (1) what can the public effectively learn from the reported data and does this information suffice to assess the NSL usage? and (2) how accessible is this data collection? On the one hand, we show that after consolidating and processing data, longitudinal trends in the usage of NSLs can be observed. For instance, we find a significant increase in NSL requests for non-US persons starting from 2015. The observed trends suggest that the current transparency mechanisms are viable safeguards against the excessive use of NSLs. On the other hand, we find that aggregating and normalizing the data is a challenging task, and one of our main contributions is producing a normalized and accessible data set and making it available for future research. The various data sources pose major challenges as they all lack a standardized and machine-readable format. Releasing transparency reports and NSLs, after their gag order was lifted, is optional for companies and only done consistently by a few. While the government data is comparatively straightforward to find, as the publishing entities are known, processing them is challenging. Indeed, the format of the reported data changed repeatedly and the publication as scanned PDFs with continuous text requires manual reviewing, parsing, and validating of the data. We even find inconsistencies within and across data sources. Overall, the laborious data collection process hinders external and internal auditing efforts and demonstrates the need for a unified and more usable dataset for NSLs. We advocate for publishing NSL statistics in a standardized and easily processable format to facilitate third-party data analysis that ensures that Congress and the public can satisfy their role of overseeing the use of NSLs by government agencies.
@@ 3.3 Company NSLs
Figure 4: This graph shows how many years it took to publish NSLs after they were issued. Most letters were only published after the USA FREEDOM Act added procedures to challenge gag orders. The gag order time appears to decrease. However, the distribution of the unpublished 99.7% of the NSLs is unknown. Moreover, the measured gag order time appears to decrease due to a bias, as letters from recent years with longer times have not been published yet. For comparison, we plot a black line to show a trend for a simulated scenario where gag order time is sampled from the uniform distribution with the constant mean value of 5 years.
However, two important caveats need to be mentioned in connection with Figure 4. First, we only see 0.3% of all NSLs, and it is unknown whether they are a representative sample for the distribution of the remaining 99.7% of NSLs. Second, data since 2013 is increasingly skewed towards shorter gag order times, as any letter that is in the “future dates” area will only be published in the future. For instance, letters issued in 2022 with a gag order time of four years will only be published in 2026. We characterize the bias and provide a toy model to estimate the true, unbiased mean gag order duration time in Appendix B.
@@ A.2 Number of NSLs per Company
Appendix B Nondisclosure duration time
B.1 Bias characterization and numerical experiments
In this section, we characterize a bias in a measured mean value of the nondisclosure duration time and suggest a simple model to estimate, under some assumptions, the true mean value. Most importantly, we show that the decreasing measured mean value, counter intuitively, does not imply that the true mean is decreasing.
We can interpret the data points $(a,b)$ on Figure 4 as an observations from some distribution $P_{observed}(a,b)$, where $a$ is a date of issue, $b$ is a nondisclosure duration of the NSL letter, $P_{observed}$ is the probability of a letter to become publicly available, $N$ is the maximum gag order duration time, and $k$ is a current time (see Figure 6).
Figure 6: This graph shows schematically how we interpret the NSL letters as observations drawn from some distribution.
Let us denote as $P_{lifted}(b=b_{1}|a=a_{1})$ a conditional probability of a gag order to be lifted after time $b_{1}$ given the NSL is issued at time $a_{1}$. To account for the selective publishing of NSL by companies, we introduce $P_{published}({published}|a=a_{1},b=b_{1})$, a conditional probability of a letter to be published given it is issued at time $a_{1}$ and the gag order is lifted at time $b_{1}$. Finally, we introduce the indicator function $\mathds{1}\{a_{1}+b_{1}\leq k\}$ to account for the fact the we only observe NSL letters published before the current time and do not observe the future letters. Then, $P_{observed}(a,b)$ is factorized as:
$P_{observed}(a,b)\propto\mathds{1}\{a_{1}+b_{1}\leq k\}\cdot P_{published}({published}|a=a_{1},b=b_{1})\cdot P_{lifted}(b=b_{1}|a=a_{1})$
The bias in the measurement and the decreasing trend on the Figure 4 stems from the indicator function $\mathds{1}\{a_{1}+b_{1}\leq k\}$.
To simplify our model, we introduce 2 assumptions:
Assumption 1.
Probability of a letter to be published is constant
$P_{published}({published}|a=a_{1},b=b_{1})=p_{pub}$
Assumption 2.
Gag order duration time is uniformly distributed
$P_{lifted}(b=b_{1}|a=a_{1})\sim U[0,N]$
Now, having defined this distribution, we can sample from it using a Gibbs sampling approach (i.e., first, sample $a$, then separately sample $b$ using Algorithm 1)
Algorithm 1 NSL sampler
$a_{1}\sim U[0,k]$
$b_{1}\sim U[0,N]$
$pub\sim Bernoulli(p_{pub})$
if $a_{1}+b_{1}<k$ and $pub$ then
     return $(a_{1},b_{1})$
else
     return $\emptyset$
end if
Now, we sample observations using Algorithm 1 and plot polynomial regression of the observations to obtain the simulated, black curve on Figure 4.
B.2 True mean estimation
With this model, we can also account for the bias and estimate the true mean gag order duration time in the following way.
Let us fix a certain value of $a=a_{1}$. In this case, the probability that sampling one NSL returns a non-empty value, i.e., that we observe the sampled NSL in our current time, is $P_{observed}$ marginalized over $b$:
$P_{observed}(a=a_{1})=\sum_{b\in[0,N]}P_{observed}(a=a_{1},b=b_{1})=\frac{k-a_{1}}{N}\cdot p_{pub}$
Given that we have $p_{pub}$ and $k$ being constant, and $a_{1}$ being fixed, this marginalized $P_{observed}$ becomes a Bernoulli distribution:
$P_{observed}(a=a_{1})=Bernoulli(\frac{k-a_{1}}{N}\cdot p_{pub})$
And the number of observed NSLs is therefore a Binomial distribution $B(n,\frac{k-a_{1}}{N}\cdot p_{pub})$ From there, we can compute its expectation by plugging $n_{a_{1}}$, the number of published NSLs at a time $a_{1}$ (the number of Bernoulli trials), which we can obtain from ASTR data:
$E[B(n_{a_{1}},\frac{k-a_{1}}{N}\cdot p_{pub})]=\frac{k-a_{1}}{N}\cdot p_{pub}\cdot n_{a_{1}}$
Finally, we know empirical value of $E[B(n_{a_{1}},\frac{k-a_{1}}{N}\cdot p_{pub})]$ from NSL data and from Figure 4, which is the number of publicly available NSL letters issued at time $a_{1}$. Let us denote it as #$\#obs_{a_{1}}$. That gives us an expression of $N$ for certain $a=a_{1}$:
$N_{a_{1}}=\frac{k-a_{1}}{\#obs_{a_{1}}}\cdot p_{pub}\cdot n_{a_{1}}$
In our tests, unfortunately, we found that this estimator does not produce plausible values of $N$, and one of the reasons might be that the actual distribution of gag order duration time $b$ is not uniform, violating our assumption 2. It is not impossible that an improved approach can provide robust mean estimation for this scenario. We hope that this toy model might be a step towards understanding what information public can infer from the National Security Letters.